# Safe Deep Policy Adaptation

### v2

@@ I INTRODUCTION
Adaptation is a crucial aspect of autonomous systems, particularly when confronted with dynamic and uncertain environments [1, 2, 3, 4, 5, 6, 7]. In these systems, robots must continually adjust their behavior and decision-making strategies to effectively respond to changing conditions. Adaptation in the context of control systems, often referred to as adaptive control [1], is a fundamental concept in control theory. Classic adaptive control approaches the adaptation problem by continually adjusting the control system parameters to maintain or improve its performance in the presence of changing dynamics, uncertainties, or disturbances [8, 9, 10]. However, the applicability of classic adaptive control methods is often constrained by their dependence on analytical representations of system dynamics and specific system classes (e.g., systems with linear parametric uncertainties). These limitations restrict their utility in more complex systems. In the field of reinforcement learning (RL) [11], policy adaptation operates in a more general setting, often without explicit system models, but it shares the core concept of adaptation [12]. Successful policy adaptation [7, 13, 14, 15, 16, 5] empowers autonomous systems to excel in a wide range of tasks and environments, making them more versatile and capable of handling real-world challenges effectively.
Though RL policy adaptation has achieved remarkable progress in autonomous systems, safety remains a critical hurdle that limits its application in real-world control tasks, as any system must ensure that its actions do not lead to harmful consequences for itself or its surroundings. Existing end-to-end safe RL methods [17, 18, 19] are hampered in real-world deployment due to the lack of safety guarantees. Hierarchical safe RL methods that provide theoretical safety guarantees [20, 21] are susceptible to breaking when confronted with modest changes in dynamics or disturbances. To the best of our knowledge, there is no prior work that jointly tackles policy adaptation and safe reinforcement learning with safety guarantees.
@@ II-A Safe Reinforcement Learning
Safe RL methods can be divided into two categories [22]: 1) end-to-end methods and 2) hierarchical methods. One representative of end-to-end methods is Lagrangian methods that solve a primal-dual optimization problem to satisfy safety constraints [17, 18]. Another branch of end-to-end methods is direct policy optimization. Constrained policy optimization (CPO) [19] extends a standard trust-region RL algorithm to CMDP, leveraging a novel bound that relates the expected cost of two policies to their average divergence. Hierarchical methods solve the constrained learning problem by adding a safety filter upon the RL policy network, projecting unsafe actions to safe actions. Safety Layer [23] solves a QP problem of cost functions with learned dynamics to safeguard unsafe actions, but it does not provide adaptation ability and safety guarantees. ISSA [20] proposes an efficient sampling algorithm for general control barrier functions to get safe actions and guarantees zero violation during agent learning, but it requires a perfect dynamics model for a one-step prediction. Our method can be categorized in the hierarchical methods but with fewer assumptions on safety certificate functions and dynamics.
@@ II-B Safe Control
The safe control community has built a rich body of literature studying how to persistently satisfy a hard safety constraint. The most widely used methods are energy-based methods including potential field methods [24], control barrier functions [25], safe set algorithms [26], and sliding mode algorithms [27]. The method most related to our work is control barrier functions. Control barrier functions give a sufficient approach to guarantee forward invariance of the safe set under the respective control law, and can be naturally used as a safety filter upon RL policy networks.
@@ II-C Adaptive Control and Policy Adaptation
Classic adaptive control algorithms often investigate the problem of stability guarantees (e.g., Lyapunov stability) and tracking error convergence (e.g., asymptotic convergence) [28, 1]. Adaptive control via online system identification is critical to deploying robotics in the real world [2]. However, inferring all the exact physics parameters can be unnecessary and difficult [7]. Recently, there has been increased interest in studying online policy adaptation of reinforcement learning for robotics [7, 13, 14, 15, 16]. These methods train RL controllers in simulation with environmental parameters (e.g., terrains and friction parameters) and then deploy the policy on real-world robots using adaptation techniques. There are also many Meta-RL works [29, 30, 31] focusing on improving the task-wise adaptation ability. Among these policy adaptation works, the most relevant work, Rapid Motor adaptation (RMA) [7], proposes a teacher-student learning framework consisting of two components: a base policy and an adaptation module. The combination of these components and the student-teacher training scheme enables the legged robots to adapt to unseen terrains. Compared to RMA, our approach has safety guarantees and better bridges the sim-to-real gap (as shown in Section VI) by fine-tuning with few-shot real-world data. To the best of our knowledge, our work is the first to investigate the policy adaptation of safe RL in dynamics.
@@ III Problem Statement
In this work, we consider a system with control-affine dynamics, which could generally represent any rigid-body robotic system [32], as:
@@
with time step $t$, state $x_{t}\in\mathbb{R}^{n}$, action $a_{t}\in\mathbb{R}^{m}$, environment configuration $e_{t}\in\mathbb{R}^{k}$, unactuated dynamics $f:\mathbb{R}^{m}\times\mathbb{R}^{k}\rightarrow\mathbb{R}^{m}$, and actuated dynamics $g:\mathbb{R}^{m}\times\mathbb{R}^{k}\rightarrow\mathbb{R}^{n\times m}$. The environment configuration $e_{t}$ (unknown in real-world deployment) represents physical properties that may change in the system, e.g., friction, payload, disturbance, etc. Different types of configuration $e_{t}$ may exert distinct influences on the system dynamics. For instance, in the case of RC Car (as shown in Figure 2), dragging a box or a heavy chair can affect both $f$ and $g$. However, certain disturbances only impact $f$, such as an additional wind force. These environment configurations $e_{t}$ play a crucial role in adaptive control. Changes in wind or dragging can cause drones to crash [33] or force cars to collide with obstacles [21], resulting in system breakdown, especially for safety-critical systems. On the other hand, $e_{t}$ also has huge impacts on control policy performance. Therefore, our objective is to design a framework that accounts for varying environment configurations, ensuring tasks are executed with a high probability of safety guarantee while achieving high task rewards. We formulate this problem as: $\max_{\pi}\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(x_{t})\right]\text{s.t. }x_{t}\in\mathcal{C}$. The safe set $\mathcal{C}$ defines a region in the state space that the robot should always stay in. We apply the standard RL [34, 7, 35] setting for solving that, say, $\pi$ is the control policy, $r\in\mathbb{R}$ denotes reward, and $\gamma$ is the discounted factor.
@@ IV-A Phase 1: Dynamics Model Learning in Simulation
With the learned encoder $\mu_{\theta_{\mu}}$, we can train an adaptive RL [^1] While we utilize an RL agent as the control policy, our proposed SafeDPA is controller-agnostic. The RL controller can be substituted with other methods, such as Model Predictive Control (MPC). policy $\pi_{\theta_{\pi}}(x_{t},\mu_{\theta}(e_{t}))$ by integrating model-free RL algorithms (e.g. PPO [34]) with the parameter-frozen environment encoder $\mu_{\theta_{\mu}}$. RL will learn the policy that maximizes expected return $\mathbb{E}_{\tau\sim\pi}\left[\Sigma_{t=0}^{\infty}\gamma^{t}r(x_{t})\right]$ under different environment latent $z_{t}$. In our early empirical experiments, we find that the training scheme in RMA [7] (i.e., jointly training the environment encoder $\mu_{\theta_{\mu}}$ with the RL agent) does not apply for dynamics learning since the latent variable $z_{t}$ learned from RL loss does not provide sufficient context for accurate dynamics prediction. Therefore, our approach largely improves the dynamics prediction by pertaining the environment encoder only with $\mathcal{L}^{\text{DYN}}_{\theta_{f},\theta_{g},\theta_{\mu}}$ to extract a more meaningful representation for dynamics learning.
@@ IV-B Phase 2: Train Adaptation Module in Simulation
As mentioned in Section III, the environment configurations $e_{t}$ are often inaccessible in the real world. Inspired by RMA [7] and UP-OSI [36], we learn an adaptation module $\phi_{\theta}(x_{t-k:t-1},a_{t-k:t-1})$ that continuously predicts the environment latent $\hat{z}_{t}$ for dynamics model and control policy, where $\{x_{t-k:t-1},a_{t-k:t-1}\}$ is a history of state-action pairs with length $k$. We predict latent $z_{t}$ rather than environment configurations $e_{t}$, since it is more generalizable and compact to consider the impact of $e_{t}$ on dynamics than its actual value [7]. We model the adaptation module using a 1-D CNN to capture spatiotemporal correlation and train it via an iterative procedure in simulation. For each iteration, we roll out an RL agent in simulation with random episodic $e_{t}$ to collect dataset $\mathcal{D}^{\text{ADAPT}}$ with tuples $((x_{t-k:t-1},a_{t-k:t-1}),e_{t-1})$. We then train $\phi_{\theta_{\phi}}$ via supervised learning with $\mathcal{D}^{\text{ADAPT}}$ to minimize loss $\mathcal{L}^{\text{ADAPT}}_{\theta_{\phi}}=\frac{1}{|\mathcal{D}^{\text{ADAPT}}|}\sum_{t}||z_{t}-\hat{z}_{t}||^{2}_{2}$, where $z_{t}=\mu_{\theta_{\mu}}(e_{t})$ and $\hat{z}_{t}=\phi_{\theta_{\phi}}(x_{t-k:t-1},a_{t-k:t-1})$.
@@ IV-D Phase 4: Safe Filter and Real-World Deployment
To achieve safe real-world deployment, we combine the adaptation module and dynamics model and leverage the control barrier function (CBF) [25] to construct a safe filter, which can guarantee the robot always stays in the safe set $\mathcal{C}$ (i.e., forward invariance). We first construct a CBF $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$, which satisfies:
@@
where $\mathcal{C}^{h}$ is a subset of the safe set $\mathcal{C}$, and $\eta$ is a scalar representing the conservativeness of CBF. If Equation 4 is satisfied for all $x_{t}\in\mathcal{C}^{h}$, then $\mathcal{C}^{h}$ and the safe set $\mathcal{C}$ are rendered forward invariant [37, 38]. We adopt affine barrier functions as [21]: $h=p^{T}x+q$, ($p\in\mathbb{R}^{n},q\in\mathbb{R}$), for the simplicity of computation. We then formulate an optimization problem (CBF-QP) satisfying Equation 4 that can be solved at each time step:
@@ V Theoretical Analysis
We further introduce another assumption regarding the Lipschitz continuity of the (neural) dynamics. These assumptions are necessary because they provide critical quantifications into the properties of dynamics and neural networks. Otherwise, we would lack the tools to rigorously analyze the reliability of these networks [39].
@@
We consider the worst-case prediction of the $h(\hat{x}_{t+1})$, which is $\min_{x_{t},a_{t}}h(\hat{x}_{t+1})$ where $\hat{x}_{t+1}=f_{\theta_{f}}(x_{t},\hat{z_{t}})+g_{\theta_{g}}(x_{t},\hat{z_{t}})a_{t}$. The prediction error of $h(\hat{x}_{t+1})$ comes from the prediction error of $\hat{x}_{t+1}$, which can be further decoupled into two parts: 1) the approximation error of $f_{\theta_{f}}$ and $g_{\theta_{g}}$; 2) the approximation error of $\phi_{\theta_{\phi}}$. As for 1) the approximation error of $f_{\theta_{f}}$ and $g_{\theta_{g}}$, supposing we have perfect regression of $z_{t}$ (i.e., $\hat{z_{t}}=z_{t}$), denote the maximum error of $h(\hat{x}_{t+1})$ under this situation as $\Delta_{1}=h(x_{t+1})-h(f_{\theta_{f}}(x_{t},\hat{z_{t}})+g_{\theta_{g}}(x_{t},\hat{z_{t}})a_{t})<\epsilon_{f}p^{T}\mathbf{1}+\epsilon_{g}\|a\|_{1}^{max}p^{T}\mathbf{1}$. As for 2) the approximation error of $\phi_{\theta_{\phi}}$, denote the maximum error of $h(\hat{x}_{t+1})$ under this situation as $\Delta_{2}=\|\hat{z_{t}}-z_{t}\|^{max}_{1}(L_{f}+L_{f_{\theta_{f}}})+\|\hat{z_{t}}-z_{t}\|^{max}_{1}\|a\|^{max}_{1}(L_{g}+L_{g_{\theta_{g}}})<\epsilon_{z}(L_{f}+L_{f_{\theta_{f}}})+\epsilon_{z}\|a\|_{1}^{max}(L_{g}+L_{g_{\theta_{g}}})$. Combining $\Delta_{1}$ and $\Delta_{2}$, by setting $\epsilon>\Delta_{1}+\Delta_{2}$ and solving $h(\hat{x}_{t+1})\geq(1-\eta)h(x_{t})-\epsilon$, we have $h(x_{t+1})\geq(1-\eta)h(x_{t})$. The forward invariance of the safe set $\mathcal{C}$ is a natural proof following  [40]. ∎
Note that the value of $\epsilon$ depends on two types of error: 1) dynamics learning error (i.e., $\epsilon_{f}$ and $\epsilon_{g}$); and 2) adaptation regression error (i.e., $\epsilon_{z}$). As these two errors decrease (e.g., with better prediction models or fine-tuning with real-world data), $\epsilon$ will decrease as well. In practice, one can statistically estimate $\epsilon$ using fine-tuning data, but quantifying exact error bounds is a challenging machine learning problem [39, 33].

### v3

@@ I INTRODUCTION
Adaptation is a crucial aspect of autonomous systems, particularly when confronted with dynamic and uncertain environments [aastrom2013adaptive, slotine1987adaptive, lavretsky2012robust, o2022neural, huang2023datt, hanover2021performance, kumar2021rma]. In these systems, robots must continually adjust their behavior and decision-making strategies to effectively respond to changing conditions. Adaptation in the context of control systems, often referred to as adaptive control [aastrom2013adaptive], is a fundamental concept in control theory. Classic adaptive control approaches the adaptation problem by continually adjusting the control system parameters to maintain or improve its performance in the presence of changing dynamics, uncertainties, or disturbances [parks1966liapunov, khosla1985parameter, ioannou1996robust]. However, the applicability of classic adaptive control methods is often constrained by their dependence on analytical representations of system dynamics and specific system classes (e.g., systems with linear parametric uncertainties). These limitations restrict their utility in more complex systems. In the field of reinforcement learning (RL) [kaelbling1996reinforcement], policy adaptation operates in a more general setting, often without explicit system models, but it shares the core concept of adaptation [zhao2020sim]. Successful policy adaptation [kumar2021rma, lee2020learning, tan2018sim, hwangbo2019learning, peng2020learning, huang2023datt] empowers autonomous systems to excel in a wide range of tasks and environments, making them more versatile and capable of handling real-world challenges effectively.
Though RL policy adaptation has achieved remarkable progress in autonomous systems, safety remains a critical hurdle that limits its application in real-world control tasks, as any system must ensure that its actions do not lead to harmful consequences for itself or its surroundings. Existing end-to-end safe RL methods [liang2018accelerated, tessler2018reward, achiam2017constrained] are hampered in real-world deployment due to the lack of safety guarantees. Hierarchical safe RL methods that provide theoretical safety guarantees [zhao2021model, cheng2019end] are susceptible to breaking when confronted with modest changes in dynamics or disturbances. To the best of our knowledge, there is no prior work that jointly tackles policy adaptation and safe reinforcement learning with safety guarantees.
@@ II-A Safe Reinforcement Learning
Safe RL methods can be divided into two categories [zhao2023state]: 1) end-to-end methods and 2) hierarchical methods. One representative of end-to-end methods is Lagrangian methods that solve a primal-dual optimization problem to satisfy safety constraints [liang2018accelerated, tessler2018reward]. Another branch of end-to-end methods is direct policy optimization. Constrained policy optimization (CPO) [achiam2017constrained] extends a standard trust-region RL algorithm to CMDP, leveraging a novel bound that relates the expected cost of two policies to their average divergence. Hierarchical methods solve the constrained learning problem by adding a safety filter upon the RL policy network, projecting unsafe actions to safe actions. Safety Layer [dalal2018safelayer] solves a QP problem of cost functions with learned dynamics to safeguard unsafe actions, but it does not provide adaptation ability and safety guarantees. ISSA [zhao2021model] proposes an efficient sampling algorithm for general control barrier functions to get safe actions and guarantees zero violation during agent learning, but it requires a perfect dynamics model for a one-step prediction. Our method can be categorized in the hierarchical methods but with fewer assumptions on safety certificate functions and dynamics.
@@ II-B Safe Control
The safe control community has built a rich body of literature studying how to persistently satisfy a hard safety constraint. The most widely used methods are energy-based methods including potential field methods [khatib1986real], control barrier functions [ames2014control], safe set algorithms [liu2014control], and sliding mode algorithms [gracia2013reactive]. The method most related to our work is control barrier functions. Control barrier functions give a sufficient approach to guarantee forward invariance of the safe set under the respective control law, and can be naturally used as a safety filter upon RL policy networks.
@@ II-C Adaptive Control and Policy Adaptation
Classic adaptive control algorithms often investigate the problem of stability guarantees (e.g., Lyapunov stability) and tracking error convergence (e.g., asymptotic convergence) [slotine1991applied, aastrom2013adaptive]. Adaptive control via online system identification is critical to deploying robotics in the real world [slotine1987adaptive]. However, inferring all the exact physics parameters can be unnecessary and difficult [kumar2021rma]. Recently, there has been increased interest in studying online policy adaptation of reinforcement learning for robotics [kumar2021rma, lee2020learning, tan2018sim, hwangbo2019learning, peng2020learning]. These methods train RL controllers in simulation with environmental parameters (e.g., terrains and friction parameters) and then deploy the policy on real-world robots using adaptation techniques. There are also many Meta-RL works [luo2021mesa, nagabandi2018learning, he2022reinforcement] focusing on improving the task-wise adaptation ability. Among these policy adaptation works, the most relevant work, Rapid Motor adaptation (RMA) [kumar2021rma], proposes a teacher-student learning framework consisting of two components: a base policy and an adaptation module. The combination of these components and the student-teacher training scheme enables the legged robots to adapt to unseen terrains. Compared to RMA, our approach has safety guarantees and better bridges the sim-to-real gap (as shown in Section VI) by fine-tuning with few-shot real-world data. To the best of our knowledge, our work is the first to investigate the policy adaptation of safe RL in dynamics.
@@ III Problem Statement
In this work, we consider a system with control-affine dynamics, which could generally represent any rigid-body robotic system [enqvist2005linear], as:
@@
with time step $t$, state $x_{t}\in\mathbb{R}^{n}$, action $a_{t}\in\mathbb{R}^{m}$, environment configuration $e_{t}\in\mathbb{R}^{k}$, unactuated dynamics $f:\mathbb{R}^{m}\times\mathbb{R}^{k}\rightarrow\mathbb{R}^{m}$, and actuated dynamics $g:\mathbb{R}^{m}\times\mathbb{R}^{k}\rightarrow\mathbb{R}^{n\times m}$. The environment configuration $e_{t}$ (unknown in real-world deployment) represents physical properties that may change in the system, e.g., friction, payload, disturbance, etc. Different types of configuration $e_{t}$ may exert distinct influences on the system dynamics. For instance, in the case of RC Car (as shown in Figure 2), dragging a box or a heavy chair can affect both $f$ and $g$. However, certain disturbances only impact $f$, such as an additional wind force. These environment configurations $e_{t}$ play a crucial role in adaptive control. Changes in wind or dragging can cause drones to crash [liu2020robust] or force cars to collide with obstacles [cheng2019end], resulting in system breakdown, especially for safety-critical systems. On the other hand, $e_{t}$ also has huge impacts on control policy performance. Therefore, our objective is to design a framework that accounts for varying environment configurations, ensuring tasks are executed with a high probability of safety guarantee while achieving high task rewards. We formulate this problem as: $\max_{\pi}\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(x_{t})\right]\text{s.t. }x_{t}\in\mathcal{C}$. The safe set $\mathcal{C}$ defines a region in the state space that the robot should always stay in. We apply the standard RL [schulman2017proximal, kumar2021rma, schulman2015trust] setting for solving that, say, $\pi$ is the control policy, $r\in\mathbb{R}$ denotes reward, and $\gamma$ is the discounted factor.
@@ IV-A Phase 1: Dynamics Model Learning in Simulation
With the learned encoder $\mu_{\theta_{\mu}}$, we can train an adaptive RL [^1] While we utilize an RL agent as the control policy, our proposed SafeDPA is controller-agnostic. The RL controller can be substituted with other methods, such as Model Predictive Control (MPC). policy $\pi_{\theta_{\pi}}(x_{t},\mu_{\theta}(e_{t}))$ by integrating model-free RL algorithms (e.g. PPO [schulman2017proximal]) with the parameter-frozen environment encoder $\mu_{\theta_{\mu}}$. RL will learn the policy that maximizes expected return $\mathbb{E}_{\tau\sim\pi}\left[\Sigma_{t=0}^{\infty}\gamma^{t}r(x_{t})\right]$ under different environment latent $z_{t}$. In our early empirical experiments, we find that the training scheme in RMA [kumar2021rma] (i.e., jointly training the environment encoder $\mu_{\theta_{\mu}}$ with the RL agent) does not apply for dynamics learning since the latent variable $z_{t}$ learned from RL loss does not provide sufficient context for accurate dynamics prediction. Therefore, our approach largely improves the dynamics prediction by pertaining the environment encoder only with $\mathcal{L}^{\text{DYN}}_{\theta_{f},\theta_{g},\theta_{\mu}}$ to extract a more meaningful representation for dynamics learning.
@@ IV-B Phase 2: Train Adaptation Module in Simulation
As mentioned in Section III, the environment configurations $e_{t}$ are often inaccessible in the real world. Inspired by RMA [kumar2021rma] and UP-OSI [yu2017preparing], we learn an adaptation module $\phi_{\theta}(x_{t-k:t-1},a_{t-k:t-1})$ that continuously predicts the environment latent $\hat{z}_{t}$ for dynamics model and control policy, where $\{x_{t-k:t-1},a_{t-k:t-1}\}$ is a history of state-action pairs with length $k$. We predict latent $z_{t}$ rather than environment configurations $e_{t}$, since it is more generalizable and compact to consider the impact of $e_{t}$ on dynamics than its actual value [kumar2021rma]. We model the adaptation module using a 1-D CNN to capture spatiotemporal correlation and train it via an iterative procedure in simulation. For each iteration, we roll out an RL agent in simulation with random episodic $e_{t}$ to collect dataset $\mathcal{D}^{\text{ADAPT}}$ with tuples $((x_{t-k:t-1},a_{t-k:t-1}),e_{t-1})$. We then train $\phi_{\theta_{\phi}}$ via supervised learning with $\mathcal{D}^{\text{ADAPT}}$ to minimize loss $\mathcal{L}^{\text{ADAPT}}_{\theta_{\phi}}=\frac{1}{|\mathcal{D}^{\text{ADAPT}}|}\sum_{t}||z_{t}-\hat{z}_{t}||^{2}_{2}$, where $z_{t}=\mu_{\theta_{\mu}}(e_{t})$ and $\hat{z}_{t}=\phi_{\theta_{\phi}}(x_{t-k:t-1},a_{t-k:t-1})$.
@@ IV-D Phase 4: Safe Filter and Real-World Deployment
To achieve safe real-world deployment, we combine the adaptation module and dynamics model and leverage the control barrier function (CBF) [ames2014control] to construct a safe filter, which can guarantee the robot always stays in the safe set $\mathcal{C}$ (i.e., forward invariance). We first construct a CBF $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$, which satisfies:
@@
where $\mathcal{C}^{h}$ is a subset of the safe set $\mathcal{C}$, and $\eta$ is a scalar representing the conservativeness of CBF. If Equation 4 is satisfied for all $x_{t}\in\mathcal{C}^{h}$, then $\mathcal{C}^{h}$ and the safe set $\mathcal{C}$ are rendered forward invariant [agrawal2017discrete, ames2016control]. We adopt affine barrier functions as [cheng2019end]: $h=p^{T}x+q$, ($p\in\mathbb{R}^{n},q\in\mathbb{R}$), for the simplicity of computation. We then formulate an optimization problem (CBF-QP) satisfying Equation 4 that can be solved at each time step:
@@ V Theoretical Analysis
We further introduce another assumption regarding the Lipschitz continuity of the (neural) dynamics. These assumptions are necessary because they provide critical quantifications into the properties of dynamics and neural networks. Otherwise, we would lack the tools to rigorously analyze the reliability of these networks [berkenkamp2017safembrl].
@@
We consider the worst-case prediction of the $h(\hat{x}_{t+1})$, which is $\min_{x_{t},a_{t}}h(\hat{x}_{t+1})$ where $\hat{x}_{t+1}=f_{\theta_{f}}(x_{t},\hat{z_{t}})+g_{\theta_{g}}(x_{t},\hat{z_{t}})a_{t}$. The prediction error of $h(\hat{x}_{t+1})$ comes from the prediction error of $\hat{x}_{t+1}$, which can be further decoupled into two parts: 1) the approximation error of $f_{\theta_{f}}$ and $g_{\theta_{g}}$; 2) the approximation error of $\phi_{\theta_{\phi}}$. As for 1) the approximation error of $f_{\theta_{f}}$ and $g_{\theta_{g}}$, supposing we have perfect regression of $z_{t}$ (i.e., $\hat{z_{t}}=z_{t}$), denote the maximum error of $h(\hat{x}_{t+1})$ under this situation as $\Delta_{1}=h(x_{t+1})-h(f_{\theta_{f}}(x_{t},\hat{z_{t}})+g_{\theta_{g}}(x_{t},\hat{z_{t}})a_{t})<\epsilon_{f}p^{T}\mathbf{1}+\epsilon_{g}\|a\|_{1}^{max}p^{T}\mathbf{1}$. As for 2) the approximation error of $\phi_{\theta_{\phi}}$, denote the maximum error of $h(\hat{x}_{t+1})$ under this situation as $\Delta_{2}=\|\hat{z_{t}}-z_{t}\|^{max}_{1}(L_{f}+L_{f_{\theta_{f}}})+\|\hat{z_{t}}-z_{t}\|^{max}_{1}\|a\|^{max}_{1}(L_{g}+L_{g_{\theta_{g}}})<\epsilon_{z}(L_{f}+L_{f_{\theta_{f}}})+\epsilon_{z}\|a\|_{1}^{max}(L_{g}+L_{g_{\theta_{g}}})$. Combining $\Delta_{1}$ and $\Delta_{2}$, by setting $\epsilon>\Delta_{1}+\Delta_{2}$ and solving $h(\hat{x}_{t+1})\geq(1-\eta)h(x_{t})-\epsilon$, we have $h(x_{t+1})\geq(1-\eta)h(x_{t})$. The forward invariance of the safe set $\mathcal{C}$ is a natural proof following  [ames2019control]. ∎
Note that the value of $\epsilon$ depends on two types of error: 1) dynamics learning error (i.e., $\epsilon_{f}$ and $\epsilon_{g}$); and 2) adaptation regression error (i.e., $\epsilon_{z}$). As these two errors decrease (e.g., with better prediction models or fine-tuning with real-world data), $\epsilon$ will decrease as well. In practice, one can statistically estimate $\epsilon$ using fine-tuning data, but quantifying exact error bounds is a challenging machine learning problem [berkenkamp2017safembrl, liu2020robust].